Problem A.1


The graph below is made of three line segments:


-1 1 2 3 4 5 6 7 8 9 10 11 12


1


2


3


4


y


x


f(x)


g(x)


h(x)


The segments correspond to the following three functions:


f(x) = x − 2, g(x) = p


4 − (x − 6)2 + 2, h(x) = x − 6


Find the total length L of the graph between x = 2 and x = 10.

1
Expert's answer
2021-11-22T15:52:11-0500
"g(x)=x-2, 2\leq x\leq 4"

"h(x)=\sqrt{4-(x-6)^2}+2, 4\leq x\leq 8"

"f(x)=x-6, 8\leq x\leq 10"



"g(x)=x-2, 2\leq x\leq 4"


"l_1=(4-2)\sqrt{2}=2\sqrt{2}"



"h(x)=\sqrt{4-(x-6)^2}+2, 4\leq x\leq 8"

Hemicircle "(x-6)^2+(y-2)^2=4, 2\leq y\leq4"


"l_2=\dfrac{1}{2}(2\pi(2)) =2\pi"

"f(x)=x-6, 8\leq x\leq 10"


"l_3=(10-8)\sqrt{2}=2\sqrt{2}"

"l=l_1+l_2+l_3"

"=2\sqrt{2}+2\pi +2\sqrt{2}"

"=4\sqrt{2}+2\pi \ (units)"

"l=4\sqrt{2}+2\pi \ (units)"